OpenStudy (anonymous):

m=-8 (7,0) make into equation

OpenStudy (anonymous):

is the answer y=-8x ?????

OpenStudy (saifoo.khan):

y - 0 = -8(x-7)

OpenStudy (saifoo.khan):

y - y1 = m(x-x1)

OpenStudy (anonymous):

y=-8x+56 is this right

OpenStudy (saifoo.khan):

\[\Huge \checkmark\]

OpenStudy (anonymous):

thanks for the help

OpenStudy (saifoo.khan):

welcome